# Can language models learn analogical reasoning? Investigating training objectives and comparisons to human performance

### v1

@@ 1 Introduction
Analogical reasoning has the potential to extend beyond just solving proportional analogies. Analogies and analogical reasoning have value regarding scientific discovery, problem solving, human creative endeavors such as metaphor generation, as well as in the field of education Boger et al. (2019); Gentner (2002); Gentner et al. (2016); Clement (1988); Gick and Holyoak (1980); Gentner and Markman (1997). In this work, we do not focus on the semantic/morphological analogies in datasets such as BATS, but on more "complex" analogies that are closer to what is used to test analogical reasoning in humans, what Ushio et al. (2021b) referred to as "Psychometric Analogy Tests". Most of the data we use has been developed precisely to test humans. Furthermore, many of these analogies will have no verbalized relation between entities, and there are likely no ways to group instances by type of relation without over-generalizing the relation. Unlike most previous studies, we want to go beyond just exploring what pre-trained language models know about analogical reasoning. Our main contributions are as follows: we propose a training objective that allows language models to learn analogies. We find that learning is effective even with a fairly small amount of data, and approaches human performance on an unseen test set constructed for testing human analogical reasoning. Lastly, we find that fine-tuning models with this training objective does not deteriorate performance on external, but related tasks with our training scheme.
@@ 2 Related Work
While a lot of work tends to focus on performance on datasets such as BATS, there exists some datasets which were developed to test humans. Ushio et al. (2021b) explored the ability of word embeddings from various transformer models as well as some pre-trained word embeddings (such as GloVe) to solve a variety of analogies originally designed to test humans in addition to the BATS and Google analogy test set. They tested three different scoring methods that aggregated across all valid analogy permutations. Their experiments found that the analogies designed for humans were much harder for all word embeddings to solve, with the better performing models’ accuracy being less than 60% on datasets designed for humans, while the highest accuracy attained on Google and BATS was 96.6% and 81.2% respectively.
Czinczoll et al. (2022) introduced a dataset comprised of scientific and metaphorical analogies, called SCAN, also testing performance on transformer based embeddings by using a natural language prompt, and filling in an ending [MASK] token. They experimented with a zero and one-shot learning setting, and additionally fine-tuned models on the BATS dataset. Overall they found that performance on the SCAN was low, and that models which had been fine-tuned on BATS lead to a decreased performance on the SCAN dataset, which the authors attribute to the types of analogies being inherently different and that datasets such as the BATS did not well represent how humans utilize analogies.
In this work, we test a novel training objective to explore whether analogical reasoning in itself is a task that can be learned, and later test to unseen analogies with non-equivalent relation types. There have been a few works in addition to Czinczoll et al. (2022) that attempted at actively learning analogies Ushio et al. (2021a); Drozd et al. (2016). Our training focuses more on similarity in relations between entities, as opposed to similarity between entities in the analogy themselves. We use this objective to fine-tune a pre-trained BERT-base model.
Additionally, there has been a lot of work in knowledge graph representation to learn embeddings for relations between entities, some of these which explicitly utilize analogical structure to learn embeddings Liu et al. (2017). However, this area of research is out of scope as we are focusing specifically on learning to represent knowledge in contextual language models.
@@ 3 Methods
All our experiments used a pre-trained BERT-base uncased model (110M parameters) [^1] https://huggingface.co/ Devlin et al. (2019) unless stated otherwise. Additionally, all our models were trained on two variants of a classification task, a binary classification and a ranking task. First, when given a proportional analogy, they learn to label it as either positive or negative. Second, for the ranking task, we present them with both a positive analogy and its negative counterpart (how these are created is described in the Datasets section below), much like a multiple choice question with two possible answers. Whichever pair the model scores as more likely an analogy is the one the model "chooses". We chose to include this task as this is how it was formulated for the human baseline we use to compare our models, and closer to how analogy tasks are often presented to humans Rogers et al. (2017); Jones et al. (2022). Unlike the binary classification task, which will rely on a cutoff chosen to differentiate positive from negative, the decision made using ranking is relative to what a specific analogy is being compared to.
We evaluate our models in three ways: we 1) evaluate the models’ ability to learn analogy classification 2) compare the models’ ability to solve analogies to a human baseline on an unseen dataset, and 3) evaluate performance on some external but related semantic similarity tasks. Statistical significance was determined using a two-sided z-test for proportions [^2] Code available at.
@@ 3.1 SBERT-modifications
Additionally, in the original paper, sentences were fed through individually. So, for example, in a sentence similarity task which rates the similarity between two individual sentences (no attending between the sentences) and the representations are determined separately. Their two individual representations are then compared. In our specific case with the $a:b::c:d$ analogy structure, the meaning of a is dependent of b and vice versa. However the meaning of c is not necessarily related to a, as they are just related to each other by their relation to their partner in their pair. Therefore we send through "a [SEP] b" separately from "c [SEP] d". This is only the case when the SBERT architecture is used. In the case that it is not, it will be explicitly stated.
@@ 3.2 Models
3.2.1 Baselines: No fine-tuning and FastText
We have three non-finetuned baselines, which we use to test the impact of our model training. The first is FastText555gensim library: https://radimrehurek.com/gensim/index.html, which was used because of its large vocabulary and ability to handle out of vocabulary words Bojanowski et al. (2017). These are the only non-contextual embeddings we test. Given an analogy pair $a:b::c:d$, the cosine similarity cos((a-b),(c-d)) was calculated. In addition to FastText, we also used two pre-trained SBERT baselines that differ only by the similarity they measure. The first similarity measure is calculated in the same way as described with FastText, which we refer to as BERT a-b non-tuned. The second BERT baseline we use calculates cos((a-c),(b-d)), which we refer to as BERT a-c non-tuned. These last two baselines will allow a more direct comparison to evaluate learning capabilities.
@@ 3.2.3 Model 2: BERT a-b
The rest of our models maintained formulating the task as a binary classification task, however instead of having a final classification head, we incorporated a novel training objective. We base our training off the idea that in a proportional analogy, the relationship between a and b can be framed as a-b Mikolov et al. (2013a). Given that the pairs that make up an analogy should have the same relation expressed between the entities, it means (a-b)=(c-d). We train our model to maximize the cosine similarity cos((a-b),(c-d)). We use a cosine embedding loss to train the model, with a margin of 0.
3.2.4 Model 3: Single Layer a-b
We also tested a model where we added a single feed forward layer on top of a pre-trained BERT model, before the pooling layer. We kept the pre-trained BERT weights frozen and trained only the additional layer.
3.2.5 Model 4: BERT a-c
Since the above experiments using cosine distance as a similarity measure do not take into account $sim((a-c),(b-d))$, we test the model using $cos((a-c),(b-d))$ as the training objective. This reordering is of interest because in these analogy pairs, a and b are semantically related to each other in some way (ex: nucleus and electron), but a and c are not (ex: nucleus and sun). This means that, in theory, a-c is a much larger distance in semantic space. Previous studies have shown that when it comes to analogies, word embeddings often fail to capture more long-distance relations Rogers et al. (2017). Long-distance relations (connecting two seemingly unrelated concepts) is of interest to the topic of computational creativity.
Table 1: Average Accuracy on Analogy Classification Task. Arrows next to BERT a-b and Single Layer a-b are labeling whether accuracy went up↑, down↓, or stayed the same→, as compared to BERT a-b non-tuned. BERT a-c is compared to BERT a-c non-tuned. Boldface indicates a statistically significant change (p<0.05) using a z-test for proportions.
@@ 3.2.5 Model 4: BERT a-c
Simple Classification BERT a-b Single Layer a-b BERT a-c
Category Overall Pos. Neg. Overall Pos. Neg. Overall Pos. Neg. Overall Pos. Neg.
Overall 0.66 0.73 0.58 0.72 0.71 0.73 0.73 0.74 0.73 0.51↑ 0.39 0.62
SAT 0.57 0.78 0.37 0.59↑ 0.64 0.53 0.60 0.71↑ 0.48 0.50↑ 0.38↓ 0.63
U2 0.57 0.66 0.48 0.58↑ 0.60 0.56 0.55↑ 0.60 0.51 0.53↑ 0.50↓ 0.57
U4 0.60 0.70 0.50 0.56 0.60 0.52 0.60 0.61 0.58 0.52↑ 0.50↓ 0.53
SCAN 0.71 0.74 0.68 0.81 0.77 0.86 0.83 0.80 0.86 0.50↑ 0.35 0.65
SCAN - Science 0.77 0.85 0.68 0.87 0.86 0.88 0.88 0.89 0.86 0.46↑ 0.29 0.63
SCAN - Metaphor 0.69 0.71 0.68 0.80 0.75 0.85 0.82 0.78 0.86 0.51↑ 0.36 0.66
@@ 3.3 Datasets
For fine-tuning, we use a combination of four datasets (described below). This data was balanced, with half of the data points being true analogies and half false. We generated one negative analogy from every positive analogy, where for every true $a:b::c:d$, a negative analogy $a:b::c^{\prime}:d^{\prime}$ was created. The choice of c’,d’ depended on the dataset and is described in their respective subsections. The data was split into ten parts (each equal to $\approx$10of the data). When creating each test set, we randomly selected positive analogies, and included their negative counterpart so that the model would not see either. Examples of analogies contained in the datasets used are in Table  7 in appendix A. Each model was trained ten times with one of the portions held out. Results shown are averaged across all ten runs.
@@ 3.3.1 SAT Dataset
The first dataset is comprised of 374 analogies taken from a portion of the SAT college entrance exam in the US (this section has since been discontinued), and has been used in several NLP publications Turney and Littman (2003); Ushio et al. (2021a, b). The original format was multiple choice, where, given a pair of words, the student had to choose another pair of words out of five provided that best formed an analogy with the given pair. Each question has one valid pair. One incorrect pair from the remaining four incorrect choices was chosen for each analogy to create negative samples, creating a total of 748 SAT analogies. One beneficial quality about these negative edges is that the questions were originally developed to be challenging to humans, therefore the incorrect option is not a pair of two random entities, but instead two entities that likely were chosen to be tricky.
@@ 3.3.2 U2 and U4
These analogies come from an English language learning website [^6] https://englishforeveryone.org/Topics/Analogies.html, used in some previous NLP analogy publications Ushio et al. (2021b); Boteanu and Chernova (2015). They were made for approximately ages 9-18, therefore comprising of a range of difficulty for humans. These questions were originally formatted as multiple choice as well, so negative instances were created in the same way as with the SAT data. The U2 dataset is a subset of the U4 dataset, so we removed all analogies that were present in the U2 from the U4.
@@ 3.3.3 Scientific and Creative Analogy dataset (SCAN)
The final dataset used in training is the Scientific and Creative Analogy dataset (SCAN), and is made up of analogies found in science, as well as commonly used metaphors in English literature Czinczoll et al. (2022). Instead of forming pairs of pairs (with 4 entities), each analogy pair is comprised of multi-entity relations. For example in the analogy comparing the solar system to an atom, solar system includes the entities, $sun,planet,gravity$, while atom includes the analogous entities $nucleus,electron,electromagnetism$. Each analogy pair provides $C(n,2)$ total analogies in the format $a:b::c:d$, where n is the number of entities in a topic. So, for example, in an analogy where each topic in a pair contains four entities, there are $C(4,2)=6$ total analogies. While analogies that make up the evaluation test were not seen during training, the entities in the pairs that make up the evaluation analogies may have been seen, given the multi-entity nature of this dataset. This allows us to test the model’s ability to learn to infer analogical relations when it is given other relations the entities do and do not have. Negative edges were created by randomly shuffling the c,d terms in the dataset. Given that the analogies were formed from combinations, random shuffling may accidentally result in a true analogy. All negative analogies were checked to make sure that they were not actually present in the positive analogy group. This created a total of 3102 analogies from this dataset - this represents almost 65% of the total data.
@@ 3.3.4 Human Baseline Comparison: Distractor Dataset
These analogies were compiled by researchers in a university psychology department, where they tested whether semantic similarity affected an adult human’s ability to correctly solve analogies Jones et al. (2022). This dataset was not used in our model training, and recently has been used to probe large language model for analogical reasoning ability Webb et al. (2023). In the original paper, the human subjects are presented with an incomplete analogy $a:b::c:?$, where they must choose between two options for the d term. There are two levels of semantic similarity the authors explored. First they test human’s abilities to solve analogies with regards to how related the c,d term is to the a,b term. Analogies are grouped into near analogies, where the a,b entities are semantically similar to the c,d entities, and far analogies, where the a,b entities are not semantically similarity to the c,d pair. Then within each of these groups, they come up with two incorrect d options, which they refer to as distractors (the incorrect choices for d). One of the incorrect d entities is more semantically similar to the c term than the true d term is to the c term, which they refer to as a high distractor salience. For example, a true analogy they use is $tarantula:spider::bee:insect$. They replace insect with hive as it is more related to bee. They measured semantic distance using LSA. The second incorrect d term that was chosen was less semantically similar (ex: replacing insect with yellow), which they refer to as low distractor salience. They also test three types of analogical relations: categorical, compositional and causal. Definitions and examples of these relations can be found in Jones et al. (2022)
Table 2: Average Accuracy on the Analogy Ranking Task. Arrows next to BERT a-b and Single Layer a-b are labeling whether accuracy went up↑, down↓, or stayed the same→, as compared to BERT a-b non-tuned. BERT a-c is compared to BERT a-c non-tuned. Boldface indicates a statistically significant change (p<0.05) using a z-test for proportions.
Baselines Fine-tuned Models
FastText BERT BERT Simple BERT Single Layer BERT
Category a-b non-tuned a-c non-tuned Classification a-b a-b a-c
OVERALL 0.81 0.69 0.46 0.54 0.84 0.87 0.50
SAT 0.87 0.63 0.47 0.52 0.82 0.87 0.49↑
U2 0.76 0.71 0.49 0.52 0.84 0.85 0.52↑
U4 0.75 0.71 0.43 0.56 0.84 0.86 0.51
SCAN 0.82 0.70 0.46 0.55 0.85 0.87 0.50
SCAN - Science 0.91 0.69 0.46 0.56 0.81 0.84 0.49↑
SCAN - Metaphor 0.80 0.70 0.46 0.55 0.86 0.88 0.50

### v4

@@ 1 Introduction
Analogical reasoning has the potential to extend beyond just solving proportional analogies. Analogies and analogical reasoning have value regarding scientific discovery, problem solving, human creative endeavors such as metaphor generation, as well as in the field of education Boger et al. (2019); Gentner (2002); Gentner et al. (2016); Clement (1988); Gick and Holyoak (1980); Gentner and Markman (1997). In this work, we do not focus on the semantic/morphological analogies in datasets such as BATS, but on more "complex" analogies that are closer to what is used to test analogical reasoning in humans, what Ushio et al. (2021b) referred to as "Psychometric Analogy Tests". Most of the data we use has been developed precisely to test humans. Furthermore, many of these analogies will have no verbalized relation between entities available, and there are likely no ways to group instances by type of relation without over-generalizing the relation. Unlike most previous studies, we want to go beyond just exploring what pretrained language models know about analogical reasoning. Our main contributions are as follows: we propose a training objective that allows language models to learn analogies. We find that learning is effective even with a fairly small amount of data, and approaches human performance on an unseen test set constructed for testing human analogical reasoning. Lastly, we find that fine-tuning models with this training objective does not deteriorate performance on external, but related tasks. All datasets are in English.
@@ 2 Related Work
While a lot of work tends to focus on performance on datasets such as BATS, there exist some datasets which were developed to test humans. Ushio et al. (2021b) explored the ability of word embeddings from various transformer models as well as some pretrained word embeddings (such as GloVe) to solve a variety of analogies originally designed to test humans in addition to the BATS and Google analogy test set. They tested three different scoring methods that aggregated across all valid analogy permutations. Their experiments found that the analogies designed for humans were much harder for all word embeddings to solve, with the better performing models’ accuracy being less than 60% on datasets designed for humans, while the highest accuracy attained on Google and BATS was 96.6% and 81.2% respectively.
Czinczoll et al. (2022) introduced a dataset composed of scientific and metaphorical analogies, called SCAN, also testing performance on transformer based embeddings by using a natural language prompt, and filling in an ending [MASK] token. They experimented with a zero and one-shot learning setting, and additionally fine-tuned models on the BATS dataset. Overall they found that performance on the SCAN was low, and that models which had been fine-tuned on BATS lead to a decreased performance on the SCAN dataset, which the authors attribute to the types of analogies being inherently different and that datasets such as the BATS did not well represent how humans utilize analogies.
In this work, we test a novel training objective to explore whether analogical reasoning in itself is a task that can be learned, and later test to unseen analogies with non-equivalent relation types. There have been a few works in addition to Czinczoll et al. (2022) that attempted at actively learning analogies or relations between words Ushio et al. (2021a); Drozd et al. (2016) using language models. Czinczoll et al. (2022) and Drozd et al. (2016), handle the analogy problem in the more common way of predicting a word to complete the analogy. Our training focuses more on similarity in relations between entities, as opposed to similarity between entities in the analogy themselves, and uses this objective to fine-tune a pretrained BERT-base model to solve analogies as a classification problem. Ushio et al. (2021a) proposed RelBERT, whose motivation and training scheme is most similar to our set up in that they are focused on relations and relational similarity, even if their training task is not formulated as an analogy problem specifically. However their methods involve prompting and their training data was much larger (SemEval 2012 Task 2) and involved only lexical relations. Our work involves an arguably simpler training scheme, with a much smaller training dataset that includes a wider variety of analogical relations.
Additionally, there has been a lot of work in knowledge graph representation to learn embeddings for relations between entities, some of these which explicitly utilize analogical structure to learn embeddings Liu et al. (2017). However, this area of research is out of scope as we are focusing specifically on learning to represent knowledge and relations in contextual language models.
@@ 3 Methods
All our experiments used a pretrained BERT-base uncased model (110M parameters) [^1] https://huggingface.co/ Devlin et al. (2019) unless stated otherwise. Models are trained on a binary classification task- when given an analogy, they must label it as either positive or negative.
We evaluate our models in three ways: 1) We evaluate the models’ ability to learn analogy classification using two variants of a classification task, the binary classification task used during training and a ranking task, where we present them with both a positive analogy and its negative counterpart (how these are created is described in the Datasets section below), much like a multiple choice question with two possible answers. Whichever pair the model scores as more likely an analogy is the one the model "chooses". We chose to include this task as this is how it was formulated for the human baseline we use to compare our models, and is closer to how analogy tasks are often presented to humans Rogers et al. (2017); Jones et al. (2022). Unlike the binary classification task, which will rely on a cutoff chosen to differentiate positive from negative, the decision made using ranking is relative to what a specific analogy is being compared to. 2) We compare the models’ ability to solve analogies to a human baseline on an unseen dataset (also using ranking as described in the previous paragraph). 3) We evaluate performance on some external but related semantic similarity tasks. Statistical significance was determined using a two-sided z-test for proportions [^2] Code available at https://github.com/idiap/analogy-learning.
@@ 3.1 SBERT-modifications
Additionally, in the original paper, sentences were fed through individually. So, for example, in a sentence similarity task which rates the similarity between two individual sentences (no attending between the sentences), the representations are determined separately. Their two individual representations are then compared. In our specific case with the $a:b::c:d$ analogy structure, the meaning of a is dependent on b and vice versa. However the meaning of c is not necessarily related to a, as they are just related to each other by their relation to their partner in their pair. Therefore we send through "a [SEP] b" separately from "c [SEP] d". This is only the case when the SBERT architecture is used. In the case that it is not, it will be explicitly stated.
@@ 3.2.2 Model 2: BERT a-b
The rest of our models maintained formulating the task as a binary classification task, however instead of having a final classification head, we incorporated a novel training objective. We base our training off the idea that in a proportional analogy, the relationship between a and b can be framed as a-b Mikolov et al. (2013a). Given that the pairs that make up an analogy should have the same relation expressed between the entities, it means (a-b)=(c-d). We train our model to maximize the cosine similarity cos((a-b),(c-d)) for positive examples and minimize for negative. We use a cosine embedding loss to train the model, with a margin of 0.
3.2.3 Model 3: BERT a-c
Since the above experiments using cosine distance as a similarity measure do not take into account $sim((a-c),(b-d))$, we test the model using $cos((a-c),(b-d))$ as the training objective. This reordering is of interest because in these analogy pairs, a and b are semantically related to each other in some way (ex: nucleus and electron), but a and c are generally not (ex: nucleus and sun). This means that, in theory, a-c is a much larger distance in semantic space. Previous studies have shown that when it comes to analogies, word embeddings often fail to capture more long-distance relations Rogers et al. (2017). Long-distance relations (connecting two seemingly unrelated concepts) is of interest to the topic of computational creativity.
@@ 3.2.3 Model 3: BERT a-c
Simple Classification BERT a-b BERT a-c
Category Overall Pos. Neg. Overall Pos. Neg. Overall Pos. Neg.
OVERALL 0.66 0.73 0.58 0.72 0.71 0.72 0.52 0.67 0.37
SAT 0.57 0.78 0.37 0.59 0.58↓ 0.65 0.53 0.67↑ 0.38
U2 0.57 0.66 0.48 0.58↑ 0.56 0.59 0.51↑ 0.62 0.40↓
U4 0.60 0.70 0.50 0.56 0.55 0.61 0.54↑ 0.71 0.36
SCAN 0.71 0.74 0.68 0.82 0.77 0.86 0.52↑ 0.67 0.37↓
SCAN - Science 0.77 0.85 0.68 0.87 0.87↑ 0.88 0.50 0.57↑ 0.43↑
SCAN - Metaphor 0.69 0.71 0.68 0.80 0.75 0.85 0.52↑ 0.69 0.35↓
Table 1: Average Accuracy on Analogy Classification Task. Arrows next to BERT a-b are labeling whether accuracy went up↑, down↓, or stayed the same→, as compared to BERT a-b non-tuned. BERT a-c is compared to BERT a-c non-tuned. Boldface indicates a statistically significant change (p<0.05) using a z-test for proportions.
3.2.4 Baselines: No fine-tuning and FastText
We have three non-finetuned baselines, which we use to test the impact of our model training. The first is FastText555gensim library: https://radimrehurek.com/gensim/index.html, which was used because of its large vocabulary and ability to handle out of vocabulary words Bojanowski et al. (2017). Given an analogy pair $a:b::c:d$, the cosine similarity cos((a-b),(c-d)) was calculated. In addition to FastText, we also used two pretrained SBERT baselines: BERT a-b non-tuned and BERT a-c non-tuned. These last two use the same cosine similarity measures as BERT a-b and BERT a-c, which will allow a more direct comparison to evaluate learning capabilities.
@@ 3.3 Datasets
For fine-tuning, we use a combination of four datasets (described below). This data was balanced, with half of the data points being true analogies and half false. We generated one negative analogy from every positive analogy, where for every true $a:b::c:d$, a negative analogy $a:b::c^{\prime}:d^{\prime}$ was created, which resultsed in a total of 4930 analogies. The choice of c’,d’ depended on the dataset and is described in their respective subsections. Examples of analogies contained in the datasets used, as well as other characteristics of the data, are in Table  7 in appendix A. The data was split into ten parts (each equal to $\approx$10of the data). When creating each test set, we randomly selected positive analogies, and included their negative counterpart so that the model would not see either. Each model was trained ten times with one of the portions held out. Results shown are averaged across all ten runs.
@@ 3.3.1 SAT Dataset
The first dataset is composed of 374 analogies taken from a portion of the SAT college entrance exam in the US (this section has since been discontinued), and has been used in several NLP publications Turney et al. (2003); Ushio et al. (2021a, b). The original format was multiple choice, where, given a pair of words, the student had to choose another pair of words out of five provided that best formed an analogy with the given pair. Each question has one valid pair. One incorrect pair from the remaining four incorrect choices was chosen for each analogy to create negative samples, creating a total of 748 SAT analogies. One beneficial quality about these negative edges is that the questions were originally developed to be challenging to humans, therefore the incorrect option is not a pair of two random entities, but instead two entities that likely were chosen to be tricky.
@@ 3.3.2 U2 and U4
These analogies come from an English language learning website [^6] https://englishforeveryone.org/Topics/Analogies.html, used in some previous NLP analogy publications Ushio et al. (2021a, b); Boteanu and Chernova (2015). They were made for approximately ages 9-18, therefore comprising of a range of difficulty for humans. These questions were originally formatted as multiple choice as well, so negative instances were created in the same way as with the SAT data. The U2 dataset is a subset of the U4 dataset, so we removed all analogies that were present in the U2 from the U4. These two datasets contributed a total of 1208 analogies.
@@ 3.3.3 Scientific and Creative Analogy dataset (SCAN)
The final dataset used in training is the Scientific and Creative Analogy dataset (SCAN), and is made up of analogies found in science, as well as commonly used metaphors in English literature Czinczoll et al. (2022). Instead of forming pairs of pairs (with 4 entities), each analogy pair is composed of multi-entity relations. For example, in the analogy comparing the solar system to an atom, solar system includes the entities, $sun,planet,gravity$, while atom includes the analogous entities $nucleus,electron,electromagnetism$. Each analogy pair provides $C(n,2)$ total analogies in the format $a:b::c:d$, where n is the number of entities in a topic. So, for example, in an analogy where each topic in a pair contains four entities, there are $C(4,2)=6$ total analogies. While analogies that make up the evaluation test were not seen during training, the entities in the pairs that make up the evaluation analogies may have been seen, given the multi-entity nature of this dataset. This allows us to test the model’s ability to learn to infer analogical relations when it is given other relations the entities do and do not have. Negative edges were created by randomly shuffling the c,d terms in the dataset. Given that the analogies were formed from combinations, random shuffling may accidentally result in a true analogy. All negative analogies were checked to make sure that they were not actually present in the positive analogy group. This created a total of 3102 analogies from this dataset - this represents about 63% of the total data.
BASELINES FINE-TUNED MODELS
FastText BERT BERT Simple BERT BERT
Category a-b non-tuned a-c non-tuned Classification a-b a-c
OVERALL 0.81 0.69 0.46 0.54 0.84 0.55
SAT 0.87 0.63 0.47 0.52 0.82 0.55
U2 0.76 0.71 0.49 0.52 0.83 0.59
U4 0.75 0.71 0.43 0.56 0.84 0.56
SCAN 0.82 0.70 0.46 0.55 0.85 0.54
SCAN - Science 0.91 0.69 0.46 0.56 0.81 0.58
SCAN - Metaphor 0.80 0.70 0.46 0.55 0.86 0.53
Table 2: Average Accuracy on the Analogy Ranking Task. Arrows next to BERT a-b are labeling whether accuracy went up↑, down↓, or stayed the same→, as compared to BERT a-b non-tuned. BERT a-c is compared to BERT a-c non-tuned. Boldface indicates a statistically significant change (p<0.05) using a z-test for proportions.
@@ 3.3.4 Human Baseline Comparison: Distractor Dataset
These analogies were compiled by researchers in a university psychology department, where they tested whether semantic similarity affected an adult human’s ability to correctly solve analogies Jones et al. (2022). This dataset was not used in our model training, and recently has been used to probe large language model for analogical reasoning ability Webb et al. (2023). In the original paper, the human subjects are presented with an incomplete analogy $a:b::c:?$, where they must choose between two options for the d term. There are two levels of semantic similarity the authors explored. First they test human’s abilities to solve analogies with regards to how related the c,d term is to the a,b term. Analogies are grouped into near analogies, where the a,b entities are semantically similar to the c,d entities, and far analogies, where the a,b entities are not semantically similar to the c,d pair. Then within each of these groups, they come up with two incorrect d options, which they refer to as distractors (the incorrect choices for d). One of the incorrect d entities is more semantically similar to the c term than the true d term is to the c term, which they refer to as a high distractor salience. For example, a true analogy they use is $tarantula:spider::bee:insect$. They replace insect with hive as it is more related to bee. They measured semantic distance using LSA. The second incorrect d term that was chosen was less semantically similar (ex: replacing insect with yellow), which they refer to as low distractor salience. They also test three types of analogical relations: categorical, compositional and causal. Definitions and examples of these relations can be found in Jones et al. (2022).